If the line 3x+ky=8 is perpendicular to the line 2x-5y=10, what is the value of k?
Found 2 solutions by Alan3354, Edwin McCravy:
Answer by Alan3354(69443)

If the line 3x+ky=8 is perpendicular to the line 2x-5y=10, what is the value of k?
------------
The slope m1 of 2x-5y=10 is 2/5.
The slope m2 of lines perpendicular is the negative inverse, m2 = -5/2
------------------
The slope of 3x+ky=8 is -3/k
-3/k = -5/2
-5k = -6
k = 6/5

Answer by Edwin McCravy(20056)